OpenStudy (anonymous):

The midpoint of UV is (5, -11). The coordinates of one endpoint are U(3, 5). Find the coordinates of endpoint V. I don't understand so can someone please help me and explain the process?

OpenStudy (polaris_s0i):

You just need to find the equation of the line and then use symmetry to figure this out. \[y = mx + b\] \[y = \frac{\Delta y}{\Delta x} x + b\] \[y = \frac{y_2 - y_1}{x_2 - x_1} x + b\] \[y = \frac{(-11) - 5}{5 - 3} x + b\] \[y = \frac{-16}{2} x + b\] \[y = -8x + b\] substitute in a point to figure out b (I'll put U in). \[5 = (-8)(3) + b\] \[5 = -24 + b\] \[b = 29\] So the equation of you're line is: \[ y = -8x + 29 \] Now the symmetry part: The midpoint is +2 in the x direction, so add another 2 to the midpoint to get V's x coordinate. \[y_3 = -8x_3 + 29\] \[y_3 = -8(7) + 29\] \[y_3 = -56 + 29\] \[y_3 = -27\] So V should be (7, -27) to check, calculate the distance from U to the midpoint. \[s = \sqrt{(\Delta y)^2 + (\Delta x)^2}\] \[s = \sqrt{(y_2 - y_1)^2 + (x_2 - x_1)^2}\] \[s = \sqrt{(-11 - 5)^2 + (5 - 3)^2}\] \[s = \sqrt{(-16)^2 + (2)^2}\] \[s = \sqrt{256 + 4}\] \[s = \sqrt{260}\] then do the same from the midpoint to V \[s = \sqrt{(\Delta y)^2 + (\Delta x)^2}\] \[s = \sqrt{(y_3 - y_2)^2 + (x_3 - x_2)^2}\] \[s = \sqrt{((-27) - (-11))^2 + (7 - 5)^2}\] \[s = \sqrt{(-16)^2 + (2)^2}\] \[s = \sqrt{256 + 4}\] \[s = \sqrt{260}\] The lengths are the same, so we are correct.